# Personalized Predictions of Glioblastoma Infiltration: Mathematical Models, Physics-Informed Neural Networks and Multimodal Scans

### v2

@@ Appendix D Additional results for patient tumors
In Fig. 14, as supplement to Fig. 5, we compare $u^{FDM}$ and $u^{PINN}$ for patient P5 using SEG data. Similar to the results in the synthetic case (Fig. 3), the predicted segmentations of $u^{PINN}$ align more closely with the actual segmentations than $u^{FDM}$ does, while the 1% contour and the predicted density is noisier than that from $u^{FDM}$. This suggests that when the model cannot fully account for the data, the PINN reduces the overall cost by reducing the data loss at the expense of the PDE loss. Accordingly, the $u^{PINN}$ metrics, including DICE scores and correlations with FET signal, tend to be better than $u^{FDM}$ metrics (see D, Table 3). This might not be desired in terms of PDE-constrained optimization, as the constraint (e.g., the PDE) is not satisfied. However, in a data-driven approach, the PDE loss can be considered as a regularization term that prevents overfitting (Balcerak et al., 2023).
@@
We show the results for patients P1-P8 using FET+SEG data in Fig.15. In Table 4, the estimated parameters P1-P8 are shown, using both SEG and FET+SEG data, and these are compared with the results from Lipková et al. (2019). We also compute various metrics to evaluate the performance of the methods, including DICE scores for both FLAIR and T1Gd segmentations, denoted as $DICE^{FLAIR}$ and $DICE^{T1Gd}$, as well as correlations with FET signals, denoted as corrFET. For results using FET+SEG data, corrFET is the correlation between the predicted FET signal and the predicted FET (Eq (9)). For results using SEG data, which does not consider the imaging model for FET signal, corrFET is the correlation between the predicted FET signal and the predicted cell density. Note that the imaging model of FET in Lipková et al. (2019) does not include the intercept A in Eq. (9).
@@
We also compare the parameters estimated from Lipková et al. (2019) with those obtained through our method. For parameters in Lipková et al. (2019), the MAP estimate of $(D,\rho,t_{\rm end})$ is used to compute $D/\rho$ and the non-dimensional $\mathcal{D}$ and $\mathcal{R}$ based on our estimated $\mathcal{D}$ and $\mathcal{R}$. We can see that $\mathcal{R}$ are similar, whereas there is a larger discrepancy in $\mathcal{D}$ values for P2, P3, P5, and P7. The $D/\rho$ ratio exhibits a larger discrepancy for P2 and P3. In terms of the metrics, the performance is similar. The observed differences could be ascribed to the Bayesian methods’ capability to explore a larger parameter space, whereas our training process tends to find parameters close to the initial guess. Despite the variation in parameters, the DICE score and correlation with the FET signal remain similar. In Lipková et al. (2019), the DICE score of T1Gd segmentation is 0, potentially due to a different definition of $y^{\rm T1Gd}$: in their method, it only comprises the enhancing region, while ours additionally includes the necrotic core. Given there is little T1Gd enhancement for this patient, the majority of the inference is dictated by their model’s attempt to fit the fact that cell density below the cut-off is not discernible outside of the T1Gd-enhancing region, resulting in a high $u_{c}^{\rm T1Gd}$.
@@
Table 4: Parameters estimated for Patients P1-P8 using FET+SEG and SEG data, compared with results from Lipková et al. ($*{}^{*}$). Parameters that are not included in the model are marked with “-”. Metrics include DICE score for both FLAIR and T1Gd segmentations, as well as correlation with the FET signal (see D).
@@
Several factors may contribute to the notably large size of the recurrent tumor. Firstly, since the post-operative tumor core region is also larger than the pre-operative tumor core region, it is possible that the resection was not complete, and the residual tumor cell keeps growing. Secondly, acute reactive changes after surgery or radiation can disrupt the fluid regulation in the brain, leading to edema (Ohmura et al., 2023). While our study assumes the edema region as part of the tumor, indicating tumor infiltration (Ohmura et al., 2023), the differentiation between peritumoral edema and glioma invasion remain a significant challenge and is an active area of research (Ohmura et al., 2023).
@@ Appendix G Correlation of density with FET signals
Two factors might contribute to the ring-like FET signals. Firstly, tumor necrosis, a common feature in GBM, results in non-viable, metabolically inactive tissue, which inherently shows reduced FET uptake (Meyer et al., 2021; la Fougère et al., 2011). Secondly, the partial volume effect might also play a role that is not captured by our image model (Soret et al., 2007) Since FET-PET spatial resolution is 4 mm, larger than the 1 mm resolution of MRI, signal averaging between active tumor regions and adjacent necrotic or non-tumor areas can lead to an underestimation of tracer uptake. These findings suggest that a more nuanced model may be required to accurately interpret FET signals, especially in the presence of heterogeneous tumor morphologies. Refining our model to account for these variations presents a promising avenue for future research.
@@ Appendix H Effect of Personalized characteristic parameters
In Table. 11, we compare the relative error of estimated parameters and the performance metrics of two cases: (1) cases S1, S2, S5, S6, which use personalized pre-training, as presented in the main text; and (2) cases $*{}^{*}$, $*{}^{*}$, $*{}^{*}$, $*{}^{*}$, which use the pre-trained network based on S7 for fine-tuning. We see that personalized pre-training recovers the parameters more accurately, and achieves better performance metrics, while using a fixed pre-trained model gives worse results. Note that S2 and S5 also have the same initial location ${\mathbf{x}_{0}}$ as S7, therefore there is no change in the geometry between pre-training and fine-tuning. We believe the suboptimal performance is mainly because the tumors exists different scales (different size and profile, as described by $\bar{L}$ and $\bar{D}/\bar{\rho}$), and the pre-trained model based on S7 can not capture the large variations in the tumor morphology. This leads to challenge in fine-tuning, as the “true” solution might be far from the initial guess, and the training might get stuck in a local minimum. Note that in this invesitagion, the pre-trained model is based on one set of characteristic parameters. Pre-training a single network for all patients with different $rD$, $rR$ and geometry will be a promising future direction.
@@
$*{}^{*}$48.90 -8.47 -9.69 -59.30 -0.01 2.28 0.60 0.58 0.90
Table 11: Percentage relative error of estimated parameters and the metrics, comparing effect of personalized pre-training (lableled S[1256]) and fixed pre-training based on pre-trained model using characteristic parameter of S7 (lableled $*{}^{*}$). Personalized characteristic parameters give better results.

### v3

@@ Appendix D Additional results for patient tumors
In Fig. 14, as supplement to Fig. 5, we compare $u^{FDM}$ and $u^{PINN}$ for patient P5 using SEG data. Similar to the results in the synthetic case (Fig. 3), the predicted segmentations of $u^{PINN}$ align more closely with the actual segmentations than $u^{FDM}$ does, while the 1% contour and the predicted density is noisier than that from $u^{FDM}$. This suggests that when the model cannot fully account for the data, the PINN reduces the overall cost by reducing the data loss at the expense of the PDE loss. Accordingly, the $u^{PINN}$ metrics, including DICE scores and correlations with FET signal, tend to be better than $u^{FDM}$ metrics (see D, Table 3). This might not be desired in terms of PDE-constrained optimization, as the constraint (e.g., the PDE) is not satisfied. However, in a data-driven approach, the PDE loss can be considered as a regularization term that prevents overfitting [3].
@@
We show the results for patients P1-P8 using FET+SEG data in Fig.15. In Table 4, the estimated parameters P1-P8 are shown, using both SEG and FET+SEG data, and these are compared with the results from Lipková et al. [43]. We also compute various metrics to evaluate the performance of the methods, including DICE scores for both FLAIR and T1Gd segmentations, denoted as $DICE^{FLAIR}$ and $DICE^{T1Gd}$, as well as correlations with FET signals, denoted as corrFET. For results using FET+SEG data, corrFET is the correlation between the predicted FET signal and the predicted FET (Eq (9)). For results using SEG data, which does not consider the imaging model for FET signal, corrFET is the correlation between the predicted FET signal and the predicted cell density. Note that the imaging model of FET in Lipková et al. (2019) does not include the intercept A in Eq. (9).
@@
We also compare the parameters estimated from Lipková et al. [43] with those obtained through our method. For parameters in Lipková et al. [43], the MAP estimate of $(D,\rho,t_{\rm end})$ is used to compute $D/\rho$ and the non-dimensional $\mathcal{D}$ and $\mathcal{R}$ based on our estimated $\mathcal{D}$ and $\mathcal{R}$. We can see that $\mathcal{R}$ are similar, whereas there is a larger discrepancy in $\mathcal{D}$ values for P2, P3, P5, and P7. The $D/\rho$ ratio exhibits a larger discrepancy for P2 and P3. In terms of the metrics, the performance is similar. The observed differences could be ascribed to the Bayesian methods’ capability to explore a larger parameter space, whereas our training process tends to find parameters close to the initial guess. Despite the variation in parameters, the DICE score and correlation with the FET signal remain similar. In Lipková et al. (2019), the DICE score of T1Gd segmentation is 0, potentially due to a different definition of $y^{\rm T1Gd}$: in their method, it only comprises the enhancing region, while ours additionally includes the necrotic core. Given there is little T1Gd enhancement for this patient, the majority of the inference is dictated by their model’s attempt to fit the fact that cell density below the cut-off is not discernible outside of the T1Gd-enhancing region, resulting in a high $u_{c}^{\rm T1Gd}$.
@@
Table 4: Parameters estimated for Patients P1-P8 using FET+SEG and SEG data, compared with results from Lipková et al. (2019)(). Parameters that are not included in the model are marked with “-”. Metrics include DICE score for both FLAIR and T1Gd segmentations, as well as correlation with the FET signal (see D).
@@
Several factors may contribute to the notably large size of the recurrent tumor. Firstly, since the post-operative tumor core region is also larger than the pre-operative tumor core region, it is possible that the resection was not complete, and the residual tumor cell keeps growing. Secondly, acute reactive changes after surgery or radiation can disrupt the fluid regulation in the brain, leading to edema [56]. While our study assumes the edema region as part of the tumor, indicating tumor infiltration [56], the differentiation between peritumoral edema and glioma invasion remain a significant challenge and is an active area of research [56].
@@ Appendix G Correlation of density with FET signals
Two factors might contribute to the ring-like FET signals. Firstly, tumor necrosis, a common feature in GBM, results in non-viable, metabolically inactive tissue, which inherently shows reduced FET uptake [55, 38]. Secondly, the partial volume effect might also play a role that is not captured by our image model [66] Since FET-PET spatial resolution is 4 mm, larger than the 1 mm resolution of MRI, signal averaging between active tumor regions and adjacent necrotic or non-tumor areas can lead to an underestimation of tracer uptake. These findings suggest that a more nuanced model may be required to accurately interpret FET signals, especially in the presence of heterogeneous tumor morphologies. Refining our model to account for these variations presents a promising avenue for future research.
@@ Appendix H Effect of Personalized characteristic parameters
In Table. 11, we compare the relative error of estimated parameters and the performance metrics of two cases: (1) cases S1, S2, S5, S6, which use personalized pre-training, as presented in the main text; and (2) cases S1, S2, S5, S6, which use the pre-trained network based on S7 for fine-tuning. We see that personalized pre-training recovers the parameters more accurately, and achieves better performance metrics, while using a fixed pre-trained model gives worse results. Note that S2 and S5 also have the same initial location ${\mathbf{x}_{0}}$ as S7, therefore there is no change in the geometry between pre-training and fine-tuning. We believe the suboptimal performance is mainly because the tumors exists different scales (different size and profile, as described by $\bar{L}$ and $\bar{D}/\bar{\rho}$), and the pre-trained model based on S7 can not capture the large variations in the tumor morphology. This leads to challenge in fine-tuning, as the “true” solution might be far from the initial guess, and the training might get stuck in a local minimum. Note that in this invesitagion, the pre-trained model is based on one set of characteristic parameters. Pre-training a single network for all patients with different $rD$, $rR$ and geometry will be a promising future direction.
@@
S6 -48.90 -8.47 -9.69 -59.30 -0.01 2.28 0.60 0.58 0.90
Table 11: Percentage relative error of estimated parameters and the metrics, comparing effect of personalized pre-training (lableled S[1256]) and fixed pre-training based on pre-trained model using characteristic parameter of S7 (lableled S[1256]). Personalized characteristic parameters give better results.